OpenStudy (anonymous):

y varies jointly as x and z, and y = 32 m, x = 6 m, and z = 8 m. What is the value of y when x = 5 m and z = 12 m?

OpenStudy (mertsj):

Can you write the joint variation equation for x y and z? Use k for the constant.

OpenStudy (anonymous):

@jim_thompson5910

OpenStudy (mertsj):

y=kxz

OpenStudy (mertsj):

Now plug in the given values and find k.

jimthompson5910 (jim_thompson5910):

does Mertsj's response make sense?

OpenStudy (anonymous):

a little, im still a little confused ..

OpenStudy (anonymous):

so it would be 32m= k *5m*12m

jimthompson5910 (jim_thompson5910):

y = kxz we're given y = 32 m, x = 6 m, and z = 8 m plug all this in to get 32 = k*6*8 now solve for k Note: I think we can ignore the m's because I think they refer to meters (eg: x = 6 m means x is 6 meters long)

OpenStudy (anonymous):

4?

jimthompson5910 (jim_thompson5910):

no

jimthompson5910 (jim_thompson5910):

32 = k*6*8 32 = k*48 k = ??

OpenStudy (anonymous):

its 4?

jimthompson5910 (jim_thompson5910):

no it's not 4 divide both sides by 48 to solve for k

OpenStudy (anonymous):

0.6?

jimthompson5910 (jim_thompson5910):

more like 0.667 but I would keep it as a fraction so k = 2/3

jimthompson5910 (jim_thompson5910):

y = kxz y = (2/3)xz y = (2/3)*5*12 plug in x = 5 and z = 12 y = ???

OpenStudy (anonymous):

my choices are 90m, 4m, 40m, 9m

jimthompson5910 (jim_thompson5910):

5*12 is what

OpenStudy (anonymous):

60

jimthompson5910 (jim_thompson5910):

now multiply that by 2 then divide that result by 3

OpenStudy (anonymous):

4

jimthompson5910 (jim_thompson5910):

close, but not quite

OpenStudy (anonymous):

40*

jimthompson5910 (jim_thompson5910):

yep

OpenStudy (anonymous):

thank you(:

jimthompson5910 (jim_thompson5910):

you're welcome